OpenStudy (anonymous):

I need to draw the graph of the line that has the given slope and that contains my given point. First Porblem Given Point: (-3, 4) Slope: m= negative 1 over 2 Second Porblem problem Point: (-1, -5) Slope m=0

OpenStudy (texaschic101):

(5,4)..x = 5 and y = 4 slope(m) = 2/3 now sub that into y = mx + b and solve for b (the y intercept) 4 = 2/3(5) + b 4 = 10/3 + b 4 - 10/3 = b 12/3 - 10/3 = b 2/3 = b so your equation is : y = 2/3x + 2/3 slope = 2/3 y intercept is (0,2/3) x intercept is : (-1,0) plot your intercepts (0,2/3) and (-1,0)...start at (-1,0) and since the slope is 2/3, go up 2 and to the right 3, then up 2 and to the right 3, and you should cross the y axis at (0,2/3) Thats the first one....do you understand ?

OpenStudy (anonymous):

I understand the first one, so does this still apply to the second and thrid one?